OpenStudy (anonymous):

solve for f f+4)/g=6

OpenStudy (anonymous):

well f + 4 = 6 (g)

OpenStudy (anonymous):

(f+4)/g=6 f+4=6/g f= 6/g -4

OpenStudy (anonymous):

then f = 6g - 4

OpenStudy (anonymous):

"g" is dividing so then this must multiply to 6 be careful

OpenStudy (anonymous):

well It was a pleasure to help you saludos desde PerĂº

OpenStudy (anonymous):

thanks